OpenStudy (anonymous):

Integrate: (2-t)(sqrt t)

OpenStudy (amistre64):

by parts ??

OpenStudy (amistre64):

u = 2-t v = integrate sqrt(t) du = -1 dv = sqrt(t)

OpenStudy (anonymous):

I'm doing a multistep problem and having trouble finding this indefinite integral

OpenStudy (anonymous):

Where do you get the -1?

OpenStudy (anonymous):

Could you right the variable expression please?

OpenStudy (raden):